OpenStudy (anonymous):

Solve the equation . 6 = x + 2/3

OpenStudy (anonymous):

[x=6-2/3]3 3x=18-2 3x=16 x=16/3

OpenStudy (anonymous):

6 = x + 2/3 -x = 2/3 - 6 = -16/3 x = 16/3